OpenStudy (rebeccapink6398):

which percent is equivalent to 5 1/20 So make the fraction into a percent?

OpenStudy (sparrow2):

multiple by 100 it will be 255%

OpenStudy (sparrow2):

i guess you have problems with percents

OpenStudy (tkhunny):

It's still parts per hundred. 5 + 1/20 = 500/100 + 5/100 = 505/100 = 505% I think sparrow interpreted your notation as \(\dfrac{51}{20}\).

OpenStudy (tkhunny):

Don't post questions without showing YOUR work.

OpenStudy (anonymous):

tkhunny is correct! Solving with fractions is always easier. Think of it as 1 = 100% but the 1 over one hundred = 1%. Make sense?

OpenStudy (rebeccapink6398):

thanks guys :D